OpenStudy (anonymous):

choose the values that would create an equation with no solution. check all that apply. 3x-6=3x-? is the awnser -6,6,3,-3 or,0

OpenStudy (texaschic101):

3x - 6 = 3x - ? 3x - 6 = 3x - (-6) 3x - 6 = 3x + 6 3x - 3x = 6 + 6 0 = 12 (no solution) 3x - 6 = 3x - 6 infinite solutions 3x - 6 = 3x - 3 3x - 3x = -3 + 6 0 = 3 (no solution) 3x - 6 = 3x - (-3) 3x - 6 = 3x + 6 3x - 3x = 6 + 6 0 = 12 (no solution) 3x - 6 = 3x - 0 3x - 3x = 6 - 0 0 = 6 (no solution)